# Toward Learning POMDPs Beyond Full-Rank Actions and State Observability

### v1

@@ 5 Experiments
Figure 3: Error bars represent standard deviation over 100 seeds. The y-axis is scaled to make convergence visible. Row 1: Estimated number of states. Row 2: Obs. matrix error relative to ground truth. Row 3: Trans. matrix error. This error is only measurable once the estimated number of states matches that of ground truth, which truncates the curves. Row 4: Total reward from planner under different sampling strategies (see Appendix C.3).
@@
Our experiments evaluate the fidelity of the learned POMDP models and explore the advantages of estimating transition and observation likelihoods. We seek to know, empirically, how quickly the learned observation matrices (and fine-grained transition matrices, if available) converge to ground truth values. We also wish to know whether the performance of the planning model is impaired by errors in the estimated similarity transform. Lastly, we evaluate whether the transition and observation likelihood estimates can be leveraged to specify a reward function to elicit desired behavior from a planner. All experiments are compared against linear PSRs and an Expectation-Maximization (EM) baseline (rabinerTutorialHidden1989; shatkayLearningGeometricallyConstrained2002) with a number of states determined by the number of components of the truncated SVD when learning a linear PSR.
@@
For our planning experiments, we verify our approach on several standard POMDPs: Tiger (kaelbling_planning_1998), T-Maze (with a truncated corridor) (bakkerReinforcementLearning2001), and Sense-Float-Reset. To allow the agent to collect an arbitrary-length string of data in all domains, we modify T-Maze to choose the next state randomly from the initial state distribution instead of terminating the sequence of interactions. Appendices B.1 and C contain details on the parameters of the learning algorithm and planner. Rewards of the original POMDPs have been learned as observations for planning.
For our reward-specification experiments, we introduce two novel domains (noisy hallway and directional hallway) whose observation and transition matrices can be fully recovered by our method. The domains share transition dynamics on a three-state ‘hallway,’ in which the actions include noisily translating left and right, choosing to deterministically stay in the current cell, or performing a reset to a uniform random state. The observation space is also shared; the agent may noisily observe whether it is on the left-end or right-end of the hallway. The domains differ in the observation distributions of the middle states. In noisy hallway, the agent noisily observes the end of the hallway in the direction of commanded movement (‘directional’ observations). In directional hallway, the agent observes left-end or right-end with probability $1/2$ (‘noisy’ observations). For more details, see Appendix 7.
@@
Convergence to true POMDP parameters. In Figure 3, our results suggest that our method successfully recovers the underlying observation models through the $L_{1}$ error of learned observation and partition-level transition likelihoods against ground truth. EM consistently converges to a local minimum and does not obtain correct observation or transition likelihoods.
@@
Planning performance with the learned model. To evaluate the performance of the planning model, we apply a standard sampling-based POMDP solver to the original ground truth POMDPs, learned PSRs, and learned POMDPs and compare the average yielded rewards. We use the sampling-based planning approach PO-UCT of silver_montecarlo_2010 with the correction described by shahNonasymptoticAnalysis2022. Ideally, planning performance should be the same across ground truth models, PSRs, and the learned partition-level POMDPs (see Appendix C.3 for discussion on rollout strategies for each model). Performance as a function of the number of action-observations collected is shown in Fig. 3, which we find to be similar across all models learned.
@@
Planning performance on specified rewards. We explore whether the likelihoods and observations yielded by our algorithm can be leveraged to direct agent behavior after the model is learned. One case, motivated by automated planning in robotics, is to direct the agent to drive a system into a set of states as determined by the states’ emitted observations by specifying a reward function (bootsClosingLearningplanning2011). After learning a POMDP, we can analyze the learned observation matrices to find the states to emit positive reward. In the past, if a PSR did not learn a reward model, then rewards were determined solely by observations (bootsClosingLearningplanning2011). Otherwise, the entire model must be relearned to estimate a reward model that depends on state (izadiPointBasedPlanning2008).
@@
Figure 4: The agent receives +1 reward for each timestep in the designated goal state (middle of hallway). ‘Obs’ refers to assigning rewards to action-observation pairs, whereas ‘state’ refers to assigning rewards to states. Error bars report standard deviation over 100 seeds.
@@
Our evaluations of this experiment are carried out on the two noisy hallway domains, where we attempt to direct the agent to drive the POMDP to the ‘middle’ hallway state with ambiguous observations. We compare the strategies of assigning rewards to observations and assigning rewards to states. In the directional domain, we assign +1+1 reward to action-observation pairs $(\texttt{left},\texttt{end-left})$ and $(\texttt{right},\texttt{end-right})$ for the former strategy, and assign +1+1 reward to the state whose maximum likelihood observations under left and right are their corresponding hallway ends for the latter. For the noisy environment, we reward the same action-observation pairs as the directional environment and also $(\texttt{left},\texttt{end-right})$ and $(\texttt{right},\texttt{end-left})$ for the former strategy, and add +1+1 reward to the state that maximizes the sum of entropy of observation distributions across all actions for the latter. The former strategy is evaluated on PSRs and POMDPs, whereas the latter is evaluated on learned POMDP models only. Performance is judged on the number of timesteps the agent spends in the desired states.
@@
Results can be found in Figure 4. In the directional domain, models that use the first strategy allow the planner to drive the middle state because it is easily identified by the observations received under left and right. The second strategy performs poorly due to slow convergence of transition matrices (see Appendix C.4). In the noisy domain, the uniform belief state and belief state that places all mass on the middle of the hallway yield the same mixture observation distribution weighted by the belief, which does not elicit the correct behavior from the planner. The planner that uses the rewards emitted from the highest-entropy state performs well after the transition matrices begin to converge. This additional flexibility highlights that learning POMDPs maintains all the advantages of PSRs and obtains the flexibility to exploit observation and transition likelihood models.
@@ 6 Related Work
Spectral methods are a common technique for learning partially-observable dynamics in the theory of RL literature (liuWhenPartially2022; zhanPACReinforcement2022). These models are sufficient to serve as black-box models for model-based RL but do not afford likelihood estimates for other computations that require general inference operations related to the latent state. Spectral methods to model learning have been applied to related settings, including linear time-invariant system identification (hoEditorialEffective1966; oymakRevisitingHo2022). Other alternate approaches for learning POMDPs have also been explored. toroicarteLearningReward2019 applies mixed-integer linear programming to learn an automaton to describe transition data. The learning problem has been framed as an automaton learning problem by angluinLearningRegular1987 to accept a particular language with strings given as data (brafmanRegularDecision2019; roncaMarkovAbstractions2022). Others have resorted to inductive logic schemes (amirLearningPartially2008; silver_learning_2021). These methods usually assume that transitions or observations made by the agent are deterministic. Other approaches have relaxed assumptions to stochastic observations but still assume deterministic transitions (deanInferringFinite1995).
Recurrent deep-learning-based architectures (wangLearningBelief2023; allenMitigatingPartial2024), can learn to make future predictions of system behavior from histories of actions and observations. Recurrent neural nets perform particularly well, unlike Transformers, which represent a fixed circuit that cannot maintain memory internally (luRethinkingTransformers2024). These recurrent models are use specialized training objectives that encourage networks to learn how to summarize histories observed by the agent (agarwal_persim_2021; allenMitigatingPartial2024) or have access to privelaged full-state observable information during training (wangLearningBelief2023). Like PSRs, the representation of the hidden state learned by these models is , and cannot readily provide likelihood models for probabilistic inference.
@@ 7 Conclusion and Future Work
We present a method that learns discrete POMDP parameters from an action-observation sequence gathered under a random exploration policy up to a partition of the state space. Our approach applies tensor decomposition methods to estimate a similarity transform to transform a PSR model to a basis where observation and transition likelihoods can be recovered. In domains where each state has a unique observation distribution aggregated across all full-rank actions, we recover the true POMDP. Otherwise, we learn the transitions between full-rank observability partitions of the state space.
In the future, we intend to improve our method to scale to larger problems, but in class and scale. Removing the restriction of learning observations from full-rank actions, or learning full transitions despite the presence of a nontrivial observability partition would be desirable. Another direction is to improve our approach to scale to larger POMDPs. Matrix-completion methods under low-rankness assumptions could help the algorithm infer missing entries in the Hankel matrix. Additional future work is to expand the theoretical foundations of our algorithm. Carefully studying our algorithm under a PAC-learning framework would contribute to our understanding of the computational complexity of learning POMDPs in general.
Acknowledgments
@@
This work was supported by ARL and the NSF Graduate Research Fellowship Program under Grants W911NF-23-2-0012 and 2141064. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DEVCOM Army Research Laboratory, the National Science Foundation, or the U.S. Government.
Reproducibility Statement: All omitted proofs to derive the theoretical claims in this paper have been included in Appendix A. Discussion regarding algorithm parameters selected (Hankel matrix size, rankness thresholds, UCT constants, etc.) and experimental domains can be found in Appendices B and C. Code, including environments and the learning algorithm, will be made available should our work be accepted for publication.
Ethics Statement: This paper discuses the derivation of a novel learning algorithm and application in toy experimental planning domains, so we do not believe there is any cause for ethical concern of our work.
Appendix A Omitted Proofs
A.1 Formalization of Assumptions
Here, we formualize the consequences of the assumtions discussed in Section 3.3.
Lemma 2.
@@ A.1 Formalization of Assumptions
Let $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{0},R,\gamma)$ be a POMDP. Let $\mathcal{D}_{n}=(a_{1},o_{1},\dots,a_{n},o_{n})$ for $n>0$, where $a_{i}\sim\operatorname{Unif}(\mathcal{A})$ for all $i$. Suppose the POMDP admits the assumptions outlined above. Furthermore, let $\hat{\mathcal{H}}$ be the ‘empirical’ Hankel matrix as computed in Eq. 6. Consider the Hankel matrix in the ‘limit of infinite data,’ where $\mathcal{H}=\lim_{n\to\infty}\hat{\mathcal{H}}$. Then $\mathcal{H}$ is the Hankel matrix of POMDP $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{\pi},R,\gamma)$ and $\operatorname{rank}(\mathcal{H})=|\mathcal{S}|$.
@@
For our proof, we rely on a fundamental result on the convergence to stationary distributions of ergodic Markov chains.
Theorem 2 (Ergodic Theorem, norrisDiscretetimeMarkov1997).
@@
Let $T$ be ergodic with stationary distribution $\pi$, and let $b_{0}$ be any initial distribution. Let $(X_{n})_{n\geq 0}$ be a Markov chain with respect to $T$ with initial distribution $b_{0}$. Then, for any bounded function $f:\mathcal{S}\to\mathbb{R}$ we have
@@
The second line applies the definition of a transition above and the proposed stationary distribution, and the third line uses the fact that $\pi$ is the stationary distribution of $T$. ∎
@@
We observe that for any $(a_{0},o_{0},s_{0},\dots,a_{k-1},o_{k-1},s_{k-1})\in\mathcal{X}^{k}$, under Lemma 3 and Theorem 2,
@@
$\hat{\mathcal{H}}_{\mathit{hist},\mathit{test}}$ $=\frac{\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},o_{i},\dots,a_{i+L},o_{i+L})=\mathit{hist}\oplus\mathit{test}}}{\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},\dots,a_{i+L},)=(a^{j_{1}},\dots,a^{j_{L}})}}$
@@
$\quad\quad\cdot\frac{\sum_{s^{m_{1}},\dots,s^{m_{L}}}\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},o_{i},s_{i}\dots,a_{i+L},o_{i+L},s_{i+L})=(a^{j_{1}},o^{k_{1}},s^{m_{1}},\dots,a^{j_{L}},o^{k_{L}},s^{m_{L}})}}{\sum_{o^{k_{1}}s^{k_{1}},\dots,o^{k_{L}}s^{k_{L}}}\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},o_{i},s_{i},\dots,a_{i+L},o_{i+L},s_{i+L})=(a^{j_{1}},o^{k_{1}},s^{m_{1}},\dots,a^{j_{L}},o^{k_{L}},s^{m_{L}})}}$
@@
$\lim_{n\to\infty}\hat{\mathcal{H}}_{\mathit{hist},\mathcal{D_{\mathrm{test}}}}$ $=\operatorname{\mathrm{P}}(o^{k_{1}},\dots,o^{k_{L}}|a^{j_{1}},\dots a^{j_{L}})$

### v2

@@ 5 Experiments
Our experiments evaluate the fidelity of the learned POMDP models and explore the advantages of estimating transition and observation likelihoods. We seek to know, empirically, how quickly the learned partition-level transition and observation matrices converge to ground truth values. We also wish to know how the performance of the planning model is impaired by estimation error from little data. Lastly, we evaluate whether the transition and observation likelihood estimates can be leveraged to specify a reward function to elicit desired behavior from a planner. All experiments are compared against linear PSRs and Expectation-Maximization (EM) [rabinerTutorialHidden1989, shatkayLearningGeometricallyConstrained2002] with a number of states determined by the number of components of the truncated SVD when learning a linear PSR.
Figure 3: The directional and noisy hallway domains. The observations el and er represent ‘end-right’ and ‘end-left’ respectively. Transitions are shared and observation distsributions differ in the middle state in the left and right actions. The right action is the mirror image of left, with the obs. likelihoods swapped in the middle state in the directional environment.
@@
For our planning experiments, we verify our approach on two standard domains, Tiger [kaelbling_planning_1998] and T-Maze (with a single corridor state) [bakkerReinforcementLearning2001], and Sense-Float-Reset. To allow the agent to collect an arbitrary-length string of data in all domains, we modify T-Maze to choose the next state randomly from the initial state distribution instead of terminating the sequence of interactions. In the supplemental material, Appendices B.1 and C contain details on the parameters of the learning algorithm and planner, including sensitivity analyses to algorithm parameters. Rewards of the original POMDPs have been learned as observations for planning. For our reward-specification experiments, we introduce two novel domains (noisy hallway and directional hallway) whose observation and transition matrices can be fully recovered by our method (see Fig. 3). For more details on all domains, see Appendix C.5.3 in the supplemental material.
Figure 4: Error bars represent standard deviation over 100 seeds. The y-axis is scaled to make convergence visible. Row 1: Estimated number of states. Row 2: Obs. matrix error relative to ground truth. Row 3: Trans. matrix error. This error is only measurable once the estimated number of states matches that of ground truth, which truncates the curves. Row 4: Total reward from planner under different models.
5.0.1 Convergence to true POMDP parameters.
@@
In Figure 4, our results suggest that our method successfully recovers the underlying observation models through the $L_{1}$ error of learned observation and partition-level transition likelihoods against ground truth. EM consistently converges to a local minimum that correctly predicts future observations but obtains incorrect POMDP matrices.
5.0.2 Planning performance with the learned model.
@@
To evaluate the performance of the learned models, we apply a standard POMDP solver to the original ground truth POMDPs, learned PSRs, and learned POMDPs. We use the sampling-based planning approach PO-UCT of Silver and Veness [silver_montecarlo_2010] with the correction described by Shah et al. [shahNonasymptoticAnalysis2022]. Ideally, average planning performance should be the same across ground truth models, PSRs, and the learned partition-level POMDPs (see Appendix C.3 for rollout strategies for each model). Fig. 4 reports that the performance across models is roughly the same.
5.0.3 Planning performance on specified rewards.
@@
We explore whether the likelihoods and observations yielded by our algorithm can be leveraged to direct agent behavior by specifying a reward function. When explicit POMDPs are available, we can analyze the learned observation matrices to directly find the states to assign positive reward. In the past, if a PSR did not learn a reward model, then rewards were determined by observations [bootsClosingLearningplanning2011]. Otherwise, the entire model must be relearned to estimate a reward model that depends on state [izadiPointBasedPlanning2008].
@@
Figure 5: The agent receives +1 reward for each timestep in the designated goal state (middle of hallway). ‘Obs’ refers to assigning rewards to action-observation pairs, whereas ‘state’ refers to assigning rewards to states. Error bars report standard deviation over 100 seeds. ‘Hacked reward’ refers to planning performance under a new reward fun.
@@
Our evaluations of this experiment are carried out on the two noisy hallway domains, where we attempt to direct the agent to drive the POMDP to the ‘middle’ hallway state with ambiguous observations (say, to collect more data for a poorly functioning perception model). We compare the strategies of assigning rewards to observations and assigning rewards to states. In the directional domain, we assign +1+1 reward to action-observation pairs $(\texttt{left},\texttt{end-left})$ and $(\texttt{right},\texttt{end-right})$ for the former strategy, and assign +1+1 reward to the state whose maximum likelihood observations under left and right are their corresponding hallway ends for the latter. For the noisy environment, we reward the same action-observation pairs as the directional environment and also $(\texttt{left},\texttt{end-right})$ and $(\texttt{right},\texttt{end-left})$ for the former strategy, and add +1+1 reward to the state that maximizes the sum of entropy of observation distributions across all actions for the latter. The former strategy is evaluated on PSRs and POMDPs, whereas the latter is evaluated on learned POMDP models only. Performance is judged on total reward gathered under the new reward function.
@@
Results can be found in Figure 5. In the directional domain, models that use the first strategy allow the planner to drive the middle state because it is easily identified by the observations received under left and right. The second strategy performs poorly with less data due to slow convergence of transition matrices (see Appendix C.4). In the noisy domain, the uniform belief state and belief state that places all mass on the middle of the hallway yield the same observation mixture distribution when weighted by the belief over states. The planner that uses observation-based rewards sees that playing the reset action or controlling the state to stay in the center to yield the same rewards, leading to unnecessary reset actions. The planner that uses the rewards emitted from the highest-entropy state correctly rewards the middle state after the transition matrices begin to converge, eliciting the correct behavior. This additional flexibility highlights that learning POMDPs maintains all the advantages of PSRs and obtains the flexibility to exploit observation and transition likelihood models.
6 On the Necessity of Observability Partitions
While learning POMDPs up to observability partitions falls short of our goal of recovering all possible POMDPs with arbitrary transition and observation matrices, the restriction of the observability partitions is in fact the best we can achieve from a single sequential trajectory. We will show that there are multiple POMDPs with different dynamics within the observability partitions that yield the same distribution over observations regardless of the actions taken, even when we assume that $\mathbf{Forw}$ and $\mathbf{Back}$ are full-rank. The principal issue is that there are multiple valid ‘forward-backward’ factorizations of the same Hankel matrix. An equivalent statement is that there is a similarity transform $P$ that converts one valid forward-backward factorization into another. Our counterexample to identifiability is framed in terms of the last point: we provide a POMDP and a similarity transformation that, when applied in the manner of Prop. 1 $P$, yields a valid $b_{0}^{\prime}$, $O^{\prime ao}$ and $T^{\prime a}$ of a different POMDP.
The counterexample is a perturbation of the sense-float-reset with three states (Fig. 1). We perturb the float transition matrix, which we associate with a new action $\texttt{float}^{\prime}$, to allow for the states to jump between the two ends of the graph. The reset, sense, and observation matrices remain unchanged.
$T^{\texttt{float}^{\prime}}=\begin{pmatrix}.4&.5&.1\\ .5&.1&.4\\ 0&.5&.5\\ \end{pmatrix}\quad P=\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&.95&.05\\ \end{pmatrix}$ (17)
It can be readily verified that after applying the matrix $P$ as a similarity transform on all POMDP matrices of the perturbed sense-float-reset problem, that we obtain another valid POMDP. Under a random exploration policy, the Hankel matrices computed from data trajectories of both POMDPs are the same, in the limit of infinite data (see Appendix A.7 in the supplemental material for proof).
Theorem 6.1
Let $\mathcal{M}=(b_{0},\mathcal{A},\mathcal{T},\mathcal{O},\mathcal{Z})$ and $\mathcal{M}^{\prime}=(b_{0}^{\prime},\mathcal{A}^{\prime},\mathcal{T}^{\prime},\mathcal{O}^{\prime},\mathcal{Z}^{\prime})$ be the two POMDPs corresponding to the POMDPs as presented above, with arbitrary initial distributions $b_{0}$ and $b_{0}^{\prime}$. Suppose the Hankel matrix $\mathcal{H}=\lim_{n\to\infty}\hat{\mathcal{H}}$, where $\hat{\mathcal{H}}$ is computed via Eq. 4 under POMDP $\mathcal{M}$, and suppose $\mathcal{H}^{\prime}$ is the analogous Hankel matrix under $\mathcal{M}^{\prime}$. Then $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are equivalent.
We can conclude that the distribution over observations is the same, regardless of the action sequences taken in either POMDP, since these observation distributions can be read from the row of either Hankel matrix row indexed by the empty history. What our counterexample suggests is that knowledge of the internal structure of a POMDP is not necessary to predict future outcomes of the system. While being able to predict future trajectories may be sufficient for planning (with goals set as a function of observations), the ambiguity of internal structure prevents us from fine-grain model manipulation such as setting goals on specific latent states for planning. Learning exact models that can be leveraged in this way can only be done when certain conditions are met.
@@ 7 Related Work
Much of the theory surrounding Hankel matrix methods has been developed to study Hidden Markov Models (HMMs), which can be viewed as POMDPs with a single action. Hankel factorization approaches to learn HMMs have been rediscovered many times [itoIdentifiabilityHidden1992, balleSpectralLearning2014, jaegerDiscretetimeDiscretevalued1998]. Since then, conditions under which the Hankel matrix represents a finite-state HMM have been investigated [andersonRealizationProblem1999, vidyasagarCompleteRealization2011]. Sample complexity and runtime complexity of these methods have also been studied in many contexts, including from sample trajectories of the HMM [huangMinimalRealization2016, sharanLearningOvercomplete2017] and from directly querying Hankel entries [mahajanLearningHidden2023].
The study of POMDP-learning has been informed by the development of Hankel methods. PSRs [littmanPredictiveRepresentations2001, singhPredictiveState2004, bootsClosingLearningplanning2011, wolfeLearningPredictive2005] were discovered separately, but share their mathematical basis with Hankel factorization approaches to Hidden Markov Models above. In the learning theory community, these spectral approaches have largely been focused on learning predictive models for model-based reinforcement learning [jinSampleEfficientReinforcement2020, chenLowerBounds2023, liuWhenPartially2022, chenPartiallyObservable2022], but not learning explicit transition and observation distributions. The study of exact HMM and POMDP recovery has largely been done using tensor decomposition methods, which require full-rankness assumptions on transition and row-stochastic observation matrices [mosselLearningNonsingular2006, huangMinimalRealization2016, azizzadenesheliReinforcementLearning2016, guoPACRL2016]. Relaxing these assumptions is the focus of this paper.
Other approaches to learning POMDPs (and HMMs) have also been explored. Unlike the stochastic case, POMDPs have been shown to be learnable when actions and observations are deterministic by matching states as equivalence classes of observed histories that produce the same outputs on future actions [angluinLearningRegular1987, rivestInferenceFinite1993, deanInferringFinite1995, brafmanRegularDecision2019, roncaMarkovAbstractions2022]. The Expectation-Maximation algorithm has been applied to both POMDPs and HMMs [rabinerTutorialHidden1989, shatkayLearningGeometricallyConstrained2002], but are often stuck in local minima. Mixed-integer programs to search through automata [toroicarteLearningReward2023] and inductive logic schemes [amirLearningPartially2008] have also been applied. An alternate view of learning deterministic discrete systems has been studied from the lens of combinatorial filters [sakcakMathematicalCharacterization2024]. Recurrent deep-learning-based architectures have also been explored [wangLearningBelief2023, allenMitigatingPartial2024, agarwalPerSimDataEfficient2021].
@@ 8 Conclusion and Future Work
We present a method that learns the parameters of a discrete POMDP from an action-observation sequence gathered under a random exploration policy up to a partition of the state space. Our approach applies tensor decomposition methods to estimate a similarity that to recover a full POMDP model under a milder set of assumptions relative to prior work. In domains where each state has a unique observation distribution aggregated across all full-rank actions, we recover the true POMDP. Otherwise, we learn the transitions between full-rank observability partitions of the state space. We also show that we can only distinguish POMDPs up to observability partitions from one sequential data trajectory.
One limitation of our work is the assumption that the forward and backward matrices are full-rank. In general, it is known that the rank of the Hankel matrix can be strictly smaller than the minimal number of states of an HMM, which is also a POMDP [jaegerDiscretetimeDiscretevalued1998, vidyasagarCompleteRealization2011]. While we have learned much from Hankel factorization algorithms for POMDP learning, the presence of this rank discrepancy suggests a broader algorithmic framework may be required to better understand the problem. Another limitation is that the algorithm only learns models from domains with few states due to the size of the requisite Hankel matrix. Prior work on HMMs characterizes computational efficiency based on the minimum size of the Hankel matrix required to ensure full-rankness [huangMinimalRealization2016, sharanLearningOvercomplete2017]. Generalizing these results to POMDPs may help us scale our algorithms to larger systems.
{credits}
8.0.1 Acknowledgements
@@
Our use of large language models is solely limited to the completion of routine coding tasks, such as experiment launching and plotting code. This work was supported by ARL and the NSF Graduate Research Fellowship Program under Grants W911NF-23-2-0012 and 2141064. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DEVCOM Army Research Laboratory, the National Science Foundation, or the U.S. Government.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
A Appendix: Omitted Proofs
A.1 Formalization of Assumptions Consequences
Lemma 2
@@ A.1 Formalization of Assumptions Consequences
Let $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{0},R,\gamma)$ be a POMDP. Suppose that the agent has collected a trajectory $\mathcal{D}_{n}=(a_{1},o_{1},\dots,a_{n},o_{n})$ for $n>0$, where $a_{i}\sim\operatorname{Unif}(\mathcal{A})$ for all $i$. Suppose the POMDP admits the assumptions outlined above. Furthermore, let $\hat{\mathcal{H}}$ be the ‘empirical’ Hankel matrix as computed in Eq. 4. Consider the Hankel matrix in the ‘limit of infinite data,’ where $\mathcal{H}=\lim_{n\to\infty}\hat{\mathcal{H}}$. Then $\mathcal{H}$ is the Hankel matrix of POMDP $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{\pi},R,\gamma)$ and $\operatorname{rank}(\mathcal{H})=|\mathcal{S}|$.
@@
Here, we prove the consequences of the assumtions discussed in Section 3.2. For our proof, we rely on a fundamental result on the convergence to stationary distributions of ergodic Markov chains.
Theorem A.1(Ergodic Theorem, [norrisDiscretetimeMarkov1997])
@@
Let $T$ be ergodic with stationary distribution $\pi$, and let $b_{0}$ be any initial distribution. Let $(X_{n})_{n\geq 0}$ be a Markov chain with respect to $T$ with initial distribution $b_{0}$. Then, for any bounded function $f:\mathcal{S}\to\operatorname{\mathbb{R}}$ we have
@@
The second line applies the definition of a transition above and the proposed stationary distribution, and the third line uses the fact that $\pi$ is the stationary distribution of $T$.
@@
We observe that for any $(a_{0},o_{0},s_{0},\dots,a_{k-1},o_{k-1},s_{k-1})\in\mathcal{X}^{k}$, under Lemma 3 and Theorem A.1,
@@
$\hat{\mathcal{H}}_{\mathit{hist},\mathit{test}}$
$=\frac{\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},o_{i},\dots,a_{i+L},o_{i+L})=\mathit{hist}\oplus\mathit{test}}}{\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},\dots,a_{i+L},)=(a^{j_{1}},\dots,a^{j_{L}})}}$
@@
$\quad\cdot\frac{\sum_{s^{m_{1}},\dots,s^{m_{L}}}\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},o_{i},s_{i}\dots,a_{i+L},o_{i+L},s_{i+L})=(a^{j_{1}},o^{k_{1}},s^{m_{1}},\dots,a^{j_{L}},o^{k_{L}},s^{m_{L}})}}{\sum_{o^{k_{1}}s^{k_{1}},\dots,o^{k_{L}}s^{k_{L}}}\sum_{i=1}^{n-L}\mathbb{I}_{(a_{i},o_{i},s_{i},\dots,a_{i+L},o_{i+L},s_{i+L})=(a^{j_{1}},o^{k_{1}},s^{m_{1}},\dots,a^{j_{L}},o^{k_{L}},s^{m_{L}})}}$
@@
$\lim_{n\to\infty}\hat{\mathcal{H}}_{\mathit{hist},\mathit{test}}$
$=\operatorname{\mathrm{P}}(o^{k_{1}},\dots,o^{k_{L}}|a^{j_{1}},\dots a^{j_{L}})$